# Self interest cumulative subtraction games

### v1

@@ 7 Related work
Our self-interest subtraction games exhibit a similar structural phenomenon. Deterministic tie-breaking is necessary to compute PSPE (well-posed recursion), but it is not always neutral. Empirically (see Appendix), we find subtraction sets for which some heaps make the supposedly indifferent player strictly better or worse under one tie-breaking rule than under another; no single rule dominates the others. As with elections and cascades, repeated local indifference can induce global divergences whenever $|S|>2$.
We now survey additional areas where tie-breaking rules have sometimes surprising consequences.
7.1 Voting
Tie-breaking in voting systems is a widely studied concern in computational social choice. Freeman, Brill, and Conitzer develop general frameworks for tie-breaking and show how the chosen method can alter both axiomatic properties and computational behavior [20]. Studies on manipulation complexity reveal that allowing flexible tie-breaking rules can increase the difficulty of manipulation [19], and that randomized tie-breaking shifts complexity trade-offs [36]. The chair’s power to break ties has been examined from a control perspective [34], demonstrating that the ability to choose among co-winners can itself serve as a computational lever. More recently, Bailey and Tovey analyze how lexicographic versus random tie-breaking affects equilibrium outcomes under strategic voting via price-of-anarchy metrics [6]. Collectively, these works support our thesis that tie-breaking conventions, though local and seemingly innocuous, shape global strategic structure.
7.2 Fair division
Tie-breaking also plays a critical role in fair division, where it can impact both perceived fairness and strategic behaviour. In the context of maximin share allocations, Procaccia and Wang show that even deterministic approximations can result in envy or inefficiency when tie-breaking is required between nearly equal outcomes [29]. Thus, local fairness does not guarantee global satisfaction. Aziz, Caragiannis, Igarashi, and Walsh study the allocation of indivisible goods and chores and show that deterministic tie-breaking can affect feasibility and tractability [4]. More generally, fair division algorithms often assume canonical or fixed tie-breaking to ensure determinism and reproducibility. Still, this determinism itself introduces systemic effects, especially in recursive or multi-agent settings. These findings align with our analysis of self-interest subtraction games: tie-breaks that appear harmless locally can induce global asymmetries.
7.3 Tournaments
In tournament theory, tie-breaking is essential to resolve outcome ambiguity in the presence of cycles and indifference. Tournament solutions refine the set of winners based on structural properties of pairwise dominance. Brandt analyzes stable sets in tournaments and introduces minimal stable sets as a refinement guaranteeing internal and external stability while respecting recursive consistency [10]. These refinements serve a role analogous to deterministic tie-breaking in dynamic games: they ensure that selections made under local symmetry yield globally coherent outcomes. Brandt, Brill, and Harrenstein offer a broader view of tournament solutions, examining axiomatic foundations and algorithmic complexity [11]. Small adjustments in refinement criteria (analogous to tie-breaking choices) can lead to drastically different solution sets.
7.4 Lexicographic games
Lexicographic objectives provide a principled way to resolve indifference in game-theoretic settings. Chatterjee et al. analyze stochastic games with lexicographic objectives and show that optimal strategies exist which respect strict preference orderings among multiple goals [12]. In a related line of work, Aghajohari, Avni, and Henzinger show that in discrete-bidding infinite-duration games, the choice of tie-breaking can influence determinacy; some rules preserve determinacy, while others do not [1]. These findings reinforce our perspective that deterministic tie-breaking conventions in cumulative games function analogously to lexicographic priority rules: locally applied, but with global strategic consequences.
7.5 Auction design
Tie-breaking plays a strategic role in auction design, especially when players are indifferent among bids or when allocation mechanisms yield multiple optimal responses. Jackson and Swinkels analyze private-value auctions and show that while equilibria can exist under various tie-breaking rules, their structure and efficiency may depend sensitively on how ties are resolved [22]. In computational settings, Chen and Peng prove that computing an approximate Bayesian Nash equilibrium in a first-price auction becomes PPAD-complete when the tie-breaking rule is part of the input [13]. Experimental evidence from Llorente-Saguer, Sheremeta, and Szech shows that favorable tie-breaking toward weaker bidders in all-pay auctions substantially alters strategic behavior and bidding intensity [33]. These studies emphasize that even seemingly minor, mechanical rules can affect equilibrium existence, efficiency, and fairness—aligning with our treatment of deterministic tie-breaking in recursive game settings.
7.6 Mechanism design
Tie-breaking is also central in mechanism design and matching markets with indifferences. Ashlagi and Nikzad [3] compare single versus multiple tie-breaking in school choice and show that relative performance depends critically on demand–supply dynamics. Ehlers [18] characterizes priority structures that admit strategy-proof tie-breaking under weak priorities, delineating when tie-breaking can be implemented without sacrificing incentives. More abstractly, Imamura and Tomoeda [21] propose the axiom ETI (equal treatment in indifference classes) for choice functions and study its compatibility with substitutability, consistency, and size monotonicity, clarifying design trade-offs. These results reinforce our view that deterministic, structured tie-breaking is a design variable rather than an afterthought.
7.7 Combinatorial auctions
Tie-breaking is a critical concern in combinatorial auctions, where multiple allocations may be equally optimal under reported valuations. In winner determination, tie-breaking can influence both revenue and allocative fairness. Conitzer and Sandholm show that even with truthful valuations, the tie-breaking rule in VCG can alter which bidder receives which bundle, raising questions about incentives and fairness [15]. In computational settings, Day and Cramton analyze core-selecting combinatorial auctions and show how deterministic tie-breaking (via quadratic payments) affects convergence and the stability of price trajectories [17]. More recently, Narosky, Spiritus, and Szymczak studied priority tie-breaking in combinatorial exchanges and showed that even slight differences in rule design can shift equilibrium outcomes and agent utilities [35]. These studies affirm that tie-breaking is a central design choice with both theoretical and practical ramifications.
7.8 Network routing
Tie-breaking also plays a critical role in graph search and network routing, particularly when multiple equally optimal paths exist. In routing protocols such as OSPF and IEEE 802.1aq, equal-cost paths are resolved using deterministic conventions (e.g., lexicographic ordering of node identifiers) to ensure consistency across distributed nodes. This mirrors heuristic search algorithms such as A where multiple nodes may share the same evaluation. Asai [2] and Corrêa et al. [16] show that tie-breaking strategies can significantly affect the performance and expansion behavior of A, and propose principled methods for prioritizing among equally valued states. More generally, Jacobs [23] develops a theoretical framework for analyzing tie-breaking in graph algorithms, highlighting the structural impact of seemingly arbitrary decisions. In dynamic settings, Bodwin and Parter [9] examine tie-breaking in shortest-path computation under edge failures, showing that consistent rules enable fast recovery and robust path selection. These results align with our approach: deterministic tie-breaking is not merely a technical detail but a structural component of recursive decision-making.
7.9 ML
Tie-breaking decisions also arise in machine learning and AI, particularly in structured prediction and search-based planning. In classification, models like decision trees and ensembles may encounter examples that fit multiple branches or classes equally well; deterministic resolution affects not only prediction stability but also interpretability. In reinforcement learning [37], when various actions have equal Q-values, consistent selection policies can influence convergence and long-run reward. Across these domains, tie-breaking converts local symmetry into globally meaningful behavior.
7.10 Richman bidding
@@
Recent work on Richman bidding combinatorial games illustrates another facet of tie-breaking. In these games, players bid to gain the right to move, paying their bid to the opponent, and ties are resolved via a marker that grants one player priority in case of equal bids. In the classical continuum setting, the theory is not sensitive to any particular tie-breaking convention [32], since equal bids occur with probability zero. In discrete settings, however, the marker becomes a strategic asset, influencing equilibrium behavior; see the survey [24]. Moreover, tie-breaking is central in [25, 26], where games generalize alternating normal play (last-move-wins): players may include the marker in their bid to counter an impending zugzwang. Switching to random or one-sided tie-breaking would lead to a different (and less rich) theory.
7.11 Extensive-form perfect-information games
@@
Tranæs [38] studies tie-breaking under indifference as a way to influence the opponent’s choices in finite extensive-form games of perfect information. A credible threat (antagonistic tie-break) selects, among tied continuations, a move that leaves the mover’s payoff unchanged while reducing the opponent’s continuation payoff, thereby inducing earlier choices favorable to the mover. Symmetrically, a credible promise (friendly tie-break) selects, under indifference, a move that increases the opponent’s continuation payoff, again steering earlier choices. These conventions are incorporated into the equilibrium concept of path-perfect subgame-perfect equilibrium: an equilibrium for which, along the path and in every subgame, there is no scope for indifference-based threats or promises to influence earlier choices profitably. Tranæs proves that every finite extensive-form game of perfect information admits a path-perfect pure subgame-perfect equilibrium.
@@ Acknowledgements
We would like to thank Eastside Preparatory School in Kirkland, WA and their Independent Curriculum / Senior Thesis program for helping support this research and surrounding work in subtraction games. The related work section has been composed together with ChatGPT in the spirit of reaching out to researchers in related areas to encourage “bridge-building” in future projects.
@@
Suppose Alice and Bob are faced with a heap of size 61, from which they may grab 4, 5, or 9 fishes. Bob has promised his Penguin clan to return with a total of 28 fishes. Alice, as the starting player, has pledged to remain friendly throughout. The question arises: should Bob maintain his friendly disposition, or should he deviate and act antagonistically? In this particular instance, it turns out that deviating benefits him. By becoming antagonistic, Bob secures one fish for immediate consumption, warming his beak, while still delivering the promised 28 fishes to his clan. If he remains friendly, however, he receives nothing for himself. Bob is puzzled, for his computational skills do not stretch beyond one heap of fish at a time: his tie-breaking rule is designed to leave him indifferent in such cases, yet here he is clearly rewarded for turning against Alice, who correspondingly loses a fish. The next day, they return to the fish pile, but the rules have changed. On their turns, they must now take 3, 8, 11, or 13 fishes, starting from a heap of size 36. Remembering the previous day’s outcome, and despite being fish-pile-wise, Bob is tempted by hunger. He once again turns antagonistic. However, this time the result is unfavorable. Bob’s clan only requires 14 fishes, knowing that Alice will start, but his tie-breaking choice leads to a total of only 13 fishes, which is one short of the requirement. His reputation within the clan is tarnished, and his beak remains unsatisfied.
@@
We will reveal discrepancies that do not occur in the two-action case and explain how they arise through backward induction. For the reader’s convenience, in the tables, we list the PSPE moves in the same columns as the outcomes.
@@
This leads to a curious effect at heap size $36$, namely Bob loses a unit by becoming antagonistic, that is, by playing in the FvA setting instead of remaining friendly in FvF. Let us justify this phenomenon, since it shows how deteministic tie-breaking becomes much more intricate in multi-action games than in the simpler two-action case. In the FvF setting, with $h=36$, Alice plays $8$, producing the outcome $(22,14)$. But in the FvA setting, while remaining friendly, she does not choose $8$, since that yields her utility $8+13=21$, whereas playing $11$ yields $11+11=22$. Conversely, in the FvF setting Alice does not choose $11$, which would lead to Bob’s utlity $13$ (the first player PSPE utility at a heap of size $25$), because by instead reducing the heap by $8$, she ensures that Bob receives $14$ units.
@@
The ruleset $S_{2}$ contains a diametral property of $S_{1}$, namely here a player can benefit strictly by unilateral deviation to antagonistic. As Lemma 7 shows, this is impossible whenever $|S|=2$. For $S_{2}$, however, such a case occurs at heap size $h=61$. By Alice removing $4$ in the FvA setting, instead of $5$ or $9$ (the PSPE moves in the FvF setting), Bob both adds one unit to his own PSPE utility, and removes one unit from Alice’s. The PSPE utilities of all options at $h=61$ are:
@@
In the FvA setting, Alice observes that all three moves give her the same utility of $32$. Acting friendly, she therefore chooses the option that maximizes Bob’s outcome, namely removing $4$, which gives him $29$ rather than $28$. By contrast, in FvF the situation is entirely different: here she can obtain $33$ by playing either $5$ or $9$, while playing $4$ yields only $32$. In FvF there is no indifference that affects Bob’s PSPE utility, and there is no reason for her to sacrifice by playing $4$. This explains the interesting discrepancies that arise in the FvA setting at $h=61$ (and many more at larger heaps). It is also interesting that whenever the deviating player gets a higher utility, the utility of the other player reduces by the same amount. This is unlike the case when the deviating player gets a lower utility, while the utility of the other player remains the same.
Purely self-interested players, as we define in this paper, aim solely to maximize their own utility, while disregarding that of their opponent. This example illustrates that a self-interested player may be strictly motivated to deviate and become antagonistic; this is a non-trivial observation while the defined outcomes do not reveal any such incentive. With only two actions, no such intrinsic motivation exists, as established earlier in this paper.

### v2

@@ 7 Discussion and related work
Our self-interest subtraction games exhibit a structural phenomenon. Deterministic tie-breaking is necessary to compute PSPE (well-posed recursion), but it is not always neutral. Empirically (see Appendix), we find subtraction sets for which some heaps make the supposedly indifferent player strictly better or worse under one tie-breaking rule than under another; no single rule dominates the others. As with elections and cascades, repeated local indifference can induce global divergences whenever $|S|>2$.
In computational social choice, a voting rule must specify what happens when two (or more) candidates are exactly tied. Deterministic conventions such as lexicographic priority versus random-candidate tie-breaking are often treated as “mere details”. However, recent work shows that the choice of tie-breaker has global consequences: it can change equilibria, manipulation incentives, and welfare. In particular, [1] analyses standard rules and proves that neither deterministic convention dominates across instances: each can make elections more or less manipulable and can change price-of-anarchy bounds. Thus, global outcome can be sensitive to how we resolve local indifferences.
In sequential social-learning models (wherein agents act one-by-one, each observing predecessors’ actions and a private signal), an informational cascade arises when agents rationally ignore their signals and follow the crowd; learning may then fail even as more signals arrive [3]. A knife-edge occurs when public information and a private signal exactly balance, and the agent is indifferent; [11] shows that the existence and properties of cascades can hinge on this local tie-breaking convention. Relaxing the usual modelling convention that an indifferent agent selects a fixed action can admit equilibria without informational cascades. Again, a local choice can have global repercussions, namely, whether society asymptotically learns the true state. For example, in a sequential learning model, a cascade can arise at knife-edge histories because an indifferent agent is forced to select a predetermined action; allowing either choice at indifference can prevent the cascade and admit equilibria in which learning continues.
@@
Let us mention an area where the choice of a tie-breaking convention has surprising theoretical consequences. In Richman bidding combinatorial games players bid to gain the right to move (in an underlying combinatorial game), paying their bid to the opponent, and ties are resolved via a marker that grants one player priority in case of equal bids. In the classical continuum setting, the theory is not sensitive to any particular tie-breaking convention [16], since equal bids occur with probability zero. In novel discrete settings, however, the marker becomes a strategic asset, influencing equilibrium behavior; see the survey [8]. Moreover, tie-breaking is central in [9, 10], where games generalize alternating normal play (last-move-wins): players may include the marker in their bid to counter an impending zugzwang. Switching to random or one-sided tie-breaking would lead to a different (and probably less rich) theory. For example, even a modest change of “not being allowed to include the marker in the bid”, induces mix strategies in equilibrium, which in itself appears to remove the standard CGT toolbox.
@@
Let us mention a related work on “extensive-form perfect-information games”, where concept such as credible threat and credible promise are close to our antagonistic and firendly settings, although in our study we have not (yet) emphazised any “threat” component. Tranæs [17] studies tie-breaking under indifference as a way to influence the opponent’s choices in finite extensive-form games of perfect information. A credible threat selects, among tied continuations, a move that leaves the mover’s payoff unchanged while reducing the opponent’s continuation payoff, and thereby inducing earlier choices favorable to the mover. Symmetrically, a credible promise selects, under indifference, a move that increases the opponent’s continuation payoff, again steering earlier choices. These conventions are incorporated into the equilibrium concept of path-perfect subgame-perfect equilibrium: an equilibrium for which, along the path and in every subgame, there is no scope for indifference-based threats or promises to influence earlier choices profitably. Tranæs proves that every finite extensive-form game of perfect information admits a path-perfect pure subgame-perfect equilibrium.
@@ Acknowledgements
We would like to thank Eastside Preparatory School in Kirkland, WA, and their Independent Curriculum/Senior Thesis program for supporting this research and related work on subtraction games. Section 7 has been composed by using also ChatGPT; we reach out to researchers in related areas to encourage overarching tie-breaking studies in future projects.
@@
Suppose Alice and Bob are faced with a heap of size 61, from which they may grab 4, 5, or 9 fishes. Bob has promised his Penguin clan to return with a total of 28 fishes. Alice, as the starting player, has pledged to remain friendly throughout. The question arises: should Bob maintain his friendly disposition, or should he deviate and act antagonistically? In this particular instance, it turns out that deviating benefits him. By becoming antagonistic, Bob secures one fish for immediate consumption, warming his beak, while still delivering the promised 28 fishes to his clan. If he remains friendly, however, he receives nothing for himself. Bob is puzzled, for his computational skills do not stretch beyond one heap of fish at a time: his tie-breaking rule is designed to leave him indifferent in such cases, yet here he is clearly rewarded for turning against Alice, who correspondingly loses a fish. The next day, they return to the fish pile, but the rules have changed. On their turns, they must now take 3, 8, 11, or 13 fishes, starting from a heap of size 36. Remembering the previous day’s outcome, and despite being fish-pile-wise, Bob is tempted by hunger. He once again turns antagonistic. However, this time the result is unfavourable. Bob’s clan only requires 14 fishes, knowing that Alice will start, but his tie-breaking choice leads to a total of only 13 fishes, which is one short of the requirement. His reputation within the clan is tarnished, and his beak remains unsatisfied.
@@
We will reveal discrepancies that do not occur in the two-action case and explain how they arise through backward induction. For the reader’s convenience, the tables list the PSPE moves in the same columns as the outcomes.
@@
This leads to a curious effect at a heap size of $36$, namely that Bob loses a unit by becoming antagonistic, that is, by playing in the FvA setting instead of remaining friendly in FvF. Let us justify this phenomenon, since it shows how deterministic tie-breaking becomes much more intricate in multi-action games than in the simpler two-action case. In the FvF setting, with $h=36$, Alice plays $8$, producing the outcome $(22,14)$. But in the FvA setting, while remaining friendly, she does not choose $8$, since that yields her utility $8+13=21$, whereas playing $11$ yields $11+11=22$. Conversely, in the FvF setting, Alice does not choose $11$, which would lead to Bob’s utility $13$ (the first player’s PSPE utility at a heap of size $25$), because by instead reducing the heap by $8$, she ensures that Bob receives $14$ units.
@@
The ruleset $S_{2}$ contains a diametral property of $S_{1}$, namely, here a player can strictly benefit by unilateral deviation to an antagonistic. As Lemma 8 shows, this is impossible whenever $|S|=2$. For $S_{2}$, however, such a case occurs at heap size $h=61$. By Alice removing $4$ in the FvA setting, instead of $5$ or $9$ (the PSPE moves in the FvF setting), Bob both adds one unit to his own PSPE utility, and removes one unit from Alice’s. The PSPE utilities of all options at $h=61$ are:
@@
In the FvA setting, Alice observes that all three moves give her the same utility of $32$. Acting friendly, she therefore chooses the option that maximizes Bob’s outcome, namely removing $4$, which gives him $29$ rather than $28$. By contrast, in FvF the situation is entirely different: here she can obtain $33$ by playing either $5$ or $9$, while playing $4$ yields only $32$. In FvF, there is no indifference that affects Bob’s PSPE utility, and there is no reason for her to sacrifice by playing $4$. This explains the interesting discrepancies that arise in the FvA setting at $h=61$ (and many more at larger heaps). It is also interesting that whenever the deviating player gets a higher utility, the utility of the other player reduces by the same amount. This is unlike the case when the deviating player gets a lower utility, while the utility of the other player remains the same.
Purely self-interested players, as we define in this paper, aim solely to maximize their own utility, while disregarding that of their opponent. This example illustrates that a self-interested player may be strictly motivated to deviate and become antagonistic; this is a non-trivial observation, while the defined outcomes do not reveal any such incentive. With only two actions, no such intrinsic motivation exists, as established earlier in this paper.